OpenStudy (anonymous):

F(x) = -x^2 –x+1 G(x) = -x^2 +x+6 Use the given functions f and g to solve for the following: (d) Solve f(x) >0 (e) Solve g(x) ≤ 0 (f) Solve f(x) > g(x) F(x) = -x^2 –x+1 G(x) = -x^2 +x+6 Use the given functions f and g to solve for the following: (d) Solve f(x) >0 (e) Solve g(x) ≤ 0 (f) Solve f(x) > g(x) @Mathematics

OpenStudy (anonymous):

first one you are going to have to set it equal zero and use the quadratic formula

OpenStudy (anonymous):

then since it is a parabola that opens down, it will be positive between the zeros. same for the second one

OpenStudy (anonymous):

how many answers are there to this ?

OpenStudy (anonymous):

not sure exactly what you mean. there will be one interval for the first problem. it contains and infinite number of numbers. lets just do it

OpenStudy (anonymous):

is this correct? -x^2-x+1>0 -x^2-x>1 x^2+x>-1 x^2+x+1/4>5/4 (x+1/2)^2>5/4 |x+1/2| > (sqrt 5) /2 1/2(1−5√)<x<12(5√−1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok- so am I missing anything to this answer?

OpenStudy (anonymous):

well sort of

OpenStudy (anonymous):

your answer is correct. but the solution contains a mistake

OpenStudy (anonymous):

1/2(1−5√)<x<1/2(5√−1)

OpenStudy (anonymous):

the correct answer you have it is \[\frac{1-\sqrt{5}}{2}<x<\frac{1+\sqrt{5}}{2}\]

OpenStudy (anonymous):

ok. Is that the entire answer? I feel like I am missing something. LOL

OpenStudy (anonymous):

that is right. for the first one

OpenStudy (anonymous):

ok- guess I am not sure of the next two

OpenStudy (anonymous):

there is a mistake in your procedure, but the answer is right. you made two mistakes and they canceled each other

OpenStudy (anonymous):

it should be \[-x^2-x+1>0\] \[-x^2-x>-1\] \[x^2+x<1\] etc.

OpenStudy (anonymous):

finally to \[|x+\frac{1}{2}|<\frac{\sqrt{5}}{2}\] an then \[\frac{-1-\sqrt{5}}{2}<x<\frac{-1+\sqrt{5}}{2}\]

OpenStudy (anonymous):

second one is easier. get \[-x^2+x+6\leq0\] \[-(x-3)(x+2)\leq 0\] \[x\leq -2\text { or } x \geq 3\]

OpenStudy (anonymous):

third one easiest of all, get \[-x^2 –x+1>-x^2 +x+6\] \[-x+1>x+6\] \[-5>2x\] \[-\frac{5}{2}>x\]

OpenStudy (anonymous):

ok- I was wondering what I did wrong the first time...I had some wierd answer- similar to what you put but a few mistakes, but this makes more sense. Thank you so much.

OpenStudy (anonymous):

this is what i had originally...(e) Since the zeros of the function are x1, 2 = −1 ± √1+24 /2 = −1 ± 5 /2= 2 and -3. The solution is, ∀ x ∈ (−∞, −3] V [2, +∞). This time the values of the zeros of the function are accepted. (f) You have to evaluate f-g>0 since f (x) − g(x) = −2x −5. The solution is ∀ x > −5/2